OpenStudy (anonymous):

Simplify : (3xy3)(2x3y) @LightYagami96

Nnesha (nnesha):

multiply remember when you multiply variables you should add their exponent \[\huge\rm x^n \times x^m = x^{n+m}\]

OpenStudy (anonymous):

3xy3(2x3y) =6x4y4

Nnesha (nnesha):

\[\huge\rm (3 \times 2) (x \times x^3) (y^3 \times y)\] now solve